mother is 18 years older than her. Her mother will be twice as old as Paula after 9 years. What is the present age of Paula?

Answer by ankor@dixie-net.com(22740)

let p = Pam's age
:
Write an equation for each statement
:
mother is 18 years older than her.
m = p + 18
Her mother will be twice as old as Paula after 9 years.
m + 9 = 2(p+9)
m + 9 = 2p + 18
m = 2p +18 - 9
m = 2p + 9
replace m with (p+18) and solve for p
What is the present age of Paula?